The population of a town increased by 2.63% per year from 2000 to 2010. The town's population at the beginning of 2000 was 74,740.
a. What was the 1-year growth factor for the town's population?
b. Write a formula for function f that models the town's population in terms of the number of years t since the beginning of 2000.
f(t)=
c. Determine the town's population at the beginning of 2010. Round your answer to the nearest whole number.

Answer by Boreal(15235)

For one year it is 1.0263
population would be Population=74740(1.0263)^t
For t=10, the population is 74740*1.0263^10=74740*1.296,
rounding only at the end, the population would be 96,894 to the nearest whole number.